# Cross-modal Cognitive Consensus guided Audio-Visual Segmentation

### v1

@@ Abstract
Audio-Visual Segmentation (AVS) aims to extract the sounding object from a video frame, which is represented by a pixel-wise segmentation mask. The pioneering work conducts this task through dense feature-level audio-visual interaction, which ignores the dimension gap between different modalities. More specifically, the audio clip could only provide a Global semantic label in each sequence, but the video frame covers multiple semantic objects across different Local regions. In this paper, we propose a Cross-modal Cognitive Consensus guided Network (C3N) to align the audio-visual semantics from the global dimension and progressively inject them into the local regions via an attention mechanism. Firstly, a Cross-modal Cognitive Consensus Inference Module (C3IM) is developed to extract a unified-modal label by integrating audio/visual classification confidence and similarities of modality-specific label embeddings. Then, we feed the unified-modal label back to the visual backbone as the explicit semantic-level guidance via a Cognitive Consensus guided Attention Module (CCAM), which highlights the local features corresponding to the interested object. Extensive experiments on the Single Sound Source Segmentation (S4) setting and Multiple Sound Source Segmentation (MS3) setting of the AVSBench dataset demonstrate the effectiveness of the proposed method, which achieves state-of-the-art performance.
@@
In [9], Zhou et al. made a new exploration toward Audio-Visual Segmentation (AVS), which focuses on extracting the sounding objects from a video frame based on their pixel-wise correspondence. This new task provides multiple modalities to facilitate more specific interesting object extraction and a dense feature-level audio-visual interaction framework is proposed to achieve this target. Despite the success in integrating multi-modal features [9], it is still challenging to well establish the semantic correspondence between audio and visual modalities due to the dimension gap. More specifically, the audio clip could only provide a Global semantic label in each sequence, but the video frame covers multiple semantic objects across different Local regions.
@@
To address the aforementioned issues, we propose a novel Cross-modal Cognitive Consensus guided Network (C3N). The schematic diagram of our method is shown in Fig 1. The C3N aligns audio and video from the global dimension via semantic-level information (i.e. classification confidence and semantic similarities) and integrates the cross-modal cognitive consensus with local visual regions through an attention mechanism. Firstly, the proposed Cross-modal Cognitive Consensus Inference Module (C3IM) feeds the extracted visual and audio features into independent heads to obtain classification confidence and convert the modality-specific labels into embedding vectors to calculate the cross-modal semantic similarities. We multiply audio/visual confidence scores and semantic similarities to infer the cross-modal cognitive consensus and extract a unified-modal label. Next, feed the unified-modal label back to hierarchical layers of the visual backbone as the global semantic-level guidance for highlighting local features corresponding to the object of interest in a channel-spatial attention manner through the parameter-free Cognitive Consensus guided Attention Module (CCAM). Finally, input the refined features into a cross-modal feature fusion module and adopt a segmentation head to integrate multi-modal features for the final segmentation.
@@ II-A Visual Object Segmentation
II-B Sound Source Localization
Recently, researchers attempt to incorporate audio to extract the visual object of interest such as the Sound Source Localization (SSL) task aims to locate the sounding object. SSL evolves from Audio-Visual Correspondence (AVC), whose pioneering work [34] fuses audio and visual features to measure audio-visual consistency. With the growth of deep learning, many SSL methods [35, 36, 37, 38, 39, 40, 41, 42] have been proposed. AVOL-Net [37] takes the middle layer output of the AVC network as the localization. Senocak et al. [36] adopts contrastive learning for audio-visual feature-level alignment. Chen et al. [40] and Lin et al. [43] improve localization performance through hard negative mining. Some recent methods [38, 39, 41, 42, 44] try to localize in multi-source scenes. Hu et al. [38, 44] propose clustering-based methods to obtain category-aware cluster centers and partition multiple objects. Qian et al. [39] tackle this task by using the class-activated map (CAM). Song et al. [41] propose a negative-free localization method achieved by visual data augmentation to mine explicit positive examples. Afouras et al. [45], Cheng et al. [46] and Tian et al. [47] propose general audio-visual frameworks, which are suitable for various downstream tasks including SSL. However, SSL generates heat maps for coarsely localizing sounding objects, which lack fine-grained descriptions.
Zhou et al. [9] propose Audio-Visual Segmentation (AVS) task that predicts pixel-level masks for sounding objects, and an AVS baseline utilizes feature-level interactions for multi-modal fusion. However, due to the lack of semantic-level guidance, it remains hard to align audio and video explicitly.
@@ II-C Label Embedding
In many cases, researchers convert label texts into embedding vectors and combine them with extracted features to enrich the framework’s semantic-level knowledge [53, 54, 55, 56, 57, 58, 59]. Zhang et al. [54] point out the importance of label information and design multiple label embedding-based models. LEAM [55] and EXAM [57] introduce interaction mechanisms to extract input-label relevance clues. GILE [56] proposes a generalized input-label strategy to strengthen the model’s performance on unseen classes. HARNN [60] incorporates recurrent layers with attention mechanism and models multi-layers dependencies. Xiong et al. [58], Cai et al. [61], and Wang et al. [62] introduce label embedding into BERT [52] or its variants to further enhance the semantic-level sensibility of the model. LTTA-LE [63] is a truncation-based method that leverages the label embedding to filter the redundant information of the long text and reduce the text length. In addition, the effectiveness of label embedding is also demonstrated in visual tasks [64, 65] and zero-shot tasks [53, 66].
@@ III Proposed Method
The overall architecture of our Cross-modal Cognitive Consensus guided Network (C3N) is illustrated in Fig. 2. The network takes an audio clip and video frames as inputs, where , and are the length, height, and width of the input video frame sequence. Firstly, we use the audio encoder and visual backbone to extract audio features and hierarchical visual features , where is the feature dimension and . Then, a Cross-modal Cognitive Consensus Inference Module (C3IM) uses pre-trained heads to compute the audio/visual classification confidence and calculate the similarities between audio and visual labels for extracting the semantic-level unified-modal label. Next, the Cognitive Consensus guided Attention Module (CCAM) highlights the local feature elements of the sounding object with the guidance of the unified-modal label. Finally, feed the refined visual features into the cross-modal feature fusion module and make predictions through the segmentation head.
III-A Cross-modal Cognitive Consensus Inference
We aim to explicitly extract global semantic-level alignment information to improve the performance of the AVS framework. Therefore, we propose a novel Cross-modal Cognitive Consensus Inference Module (C3IM) to exploit semantic-level audio/visual classification confidence and labels to infer the audio-visual cognitive consensus.
@@ III-A Cross-modal Cognitive Consensus Inference
The visual backbone is initialized with weights pre-trained on ImageNet [67], a large-scale image classification dataset totaling 1000 classes. And the audio encoder is pre-trained on AudioSet [68] with 527 audio classes. Such a diversity of classes applies to the semantic description of audio/visual objects in most scenarios. In practice, we freeze the parameters of the audio and visual classification heads. Firstly, as shown in Fig. 3, for every-second audio feature and the highest level visual feature , we compute the corresponding classification confidence scores:
@@
where and are pre-trained weights of the classification heads. and are classification confidence scores, where and are confidence of the audio class and visual class.
@@
Then, to bridge the semantic gap between distinct modalities, we measure the similarities of audio and visual labels and construct a semantic similarity matrix. The audio labels and visual labels are denoted as and respectively. We convert the audio and visual labels to a unified label embedding space by means of SpaCy, an advanced library for natural language processing to calculate similarities between label words or phrases. The audio and visual label embeddings are represented as and respectively, where the embedding dimension is 300. The similarity between the audio and visual labels is defined by the following equation:
@@
where denotes the Frobenius norm, and , are the row and column indexes of the semantic similarity matrix .
@@
where and are balance coefficients, which are set to 0.1. Values in the re-weighted matrix can also consider as cross-modal cognitive consensus degrees. With the inferred cognitive consensus, semantically relevant audio-visual objects can be identified to guide the following segmentation.
III-B Cognitive Consensus guided Attention
@@
After the cross-modal cognitive consensus inference, we feed the semantically consistent unified-modal label to the hierarchical layers of the visual backbone. In addition, we propose a parameter-free Cognitive Consensus guided Attention Module (CCAM), which injects class-activated weights into visual features to highlight the local feature elements corresponding to the sounding object.
@@
We search for the maximum value of and obtain its row and column indexes and . Specifically, and refer to the index of the audio and visual classes with the highest cognitive consensus score. Inspired by the ideology of Grad-CAM [69], we calculate the loss for the visual class and perform the backpropagation to compute the corresponding gradients. Different from Grad-CAM, which computes class-activated heat maps, we aim to obtain the weights corresponding to the unified-modal label with the highest cognitive consensus. This process can be defined as:
@@
where denotes classification logit of the visual class, denotes feature map activations, denotes the index of the channel. In practice, we take the last layer of each stage as the activation layer and the class-activated weights of stage is denoted as .
@@
To integrate the class-activated weights with the initial visual features for highlighting the local feature elements corresponding to high cognitive consensus, we propose a parameter-free CCAM, whose structure is shown in the bottom right of Fig 2. Without any learnable parameters, CCAM performs channel-spatial attention to the initial visual features and output refined features with the guidance of inferred global semantic-level unified-modal label, which can be expressed as:
@@
where denotes sigmoid activation function, denotes element-wise multiplication with boardcast mechanism. By means of the channel-spatial attention mechanism, CCAM incorporates semantic-level cognitive consensus weights with feature-level visual representations to guide the model focus on the semantically consistent visual object.
@@ III-C Cross-modal Feature Fusion
After getting the visual feature , feed it with the audio feature into a cross-modal Non-local block [70, 9], which utilizes 3-D convolutional layers to model the feature-level temporal-spatial dependencies and captures the inter-modal correlations. We first map and repeat into . Then, the cross-modal feature-level interaction process at stage can be formulated as follows:
@@ III-D Segmentation Head
To optimize the parameters of the proposed model. We adopt the binary cross entropy (BCE) loss as the loss function:
@@ IV-A1 Dataset
The proposed method is evaluated on the AVSbench [9] dataset for audio-visual segmentation. To the best of our knowledge, there is only one open-source dataset (AVSBench) in the AVS field. Videos from AVSBench are collected from YouTube and each video is cropped to 5 seconds. Video frames at the end of each second are extracted. Depending on the number of sounding objects in the video, the AVSBench dataset is divided into a Single-Source subset and a Multi-Source subset. The Single-Source subset contains 4,392 videos from 23 categories. Note that only the first sampled frame is annotated for videos in the train split, while all sampled frames of the val/test split are annotated, totaling 10,852 annotations. The Multi-Sources subset contains 424 videos and 2,120 annotated pixel-level segmentation masks. All sampled frames of the train/val/test split have annotation since the sounding objects may change over time. As with the pioneering work [9], the following experiments are carried out under the semi-supervised Single Sound Source Segmentation (S4) setting and the fully supervised Multiple Sound Source Segmentation (MS3) setting.

### v4

@@ Abstract
Audio-Visual Segmentation (AVS) aims to extract the sounding object from a video frame, which is represented by a pixel-wise segmentation mask for application scenarios such as multi-modal video editing, augmented reality, and intelligent robot systems. The pioneering work conducts this task through dense feature-level audio-visual interaction, which ignores the dimension gap between different modalities. More specifically, the audio clip could only provide a Global semantic label in each sequence, but the video frame covers multiple semantic objects across different Local regions, which leads to mislocalization of the representationally similar but semantically different object. In this paper, we propose a Cross-modal Cognitive Consensus guided Network (C3N) to align the audio-visual semantics from the global dimension and progressively inject them into the local regions via an attention mechanism. Firstly, a Cross-modal Cognitive Consensus Inference Module (C3IM) is developed to extract a unified-modal label by integrating audio/visual classification confidence and similarities of modality-agnostic label embeddings. Then, we feed the unified-modal label back to the visual backbone as the explicit semantic-level guidance via a Cognitive Consensus guided Attention Module (CCAM), which highlights the local features corresponding to the interested object. Extensive experiments on the Single Sound Source Segmentation (S4) setting and Multiple Sound Source Segmentation (MS3) setting of the AVSBench dataset demonstrate the effectiveness of the proposed method, which achieves state-of-the-art performance. Code will be available at https://github.com/ZhaofengSHI/AVS-C3N once accepted.
@@
In recent years, many outstanding cross-modal learning methods [9, 10, 11, 12, 13] have been developed to construct cross-modal consensus. In [14], Zhou et al. made a new exploration toward Audio-Visual Segmentation (AVS), which focuses on extracting the sounding objects from a video frame based on their pixel-wise correspondence with multiple application scenarios such as extracting the sounding objects in videos and customizing the background for multi-modal video editing [15, 16, 17], identifying and emphasizing sounding objects for augmented reality [18, 19], navigating towards the sounding object in the scene for intelligent robot systems [20, 21, 22], and so on. This new task provides multiple modalities to facilitate more specific interesting object extraction and a dense feature-level audio-visual interaction framework is proposed to achieve this target. Despite the success in integrating multi-modal features [14, 23, 24, 25], it is still challenging to establish the semantic correspondence between audio and visual modalities due to the dimension gap. More specifically, the audio clip could only provide a Global semantic label in each sequence, but the video frame covers multiple semantic objects across different Local regions. Such Global-Local dimension gap is difficult to fill by the feature-level interaction and leads to incorrect localization due to representationally similar but semantically different objects (e.g. a normal van and an ambulance) in a frame. Some methods [26, 27] aim to utilize audio and visual semantics to filter the mask of the sounding object among the generated potential masks via a two-stage strategy, whereas the two-stage frameworks are inconvenient to deploy and the aligned semantics are unable to directly interact with representations of the visual branch to improve the generated mask proposals.
@@
To overcome the aforementioned issues, we propose a novel Cross-modal Cognitive Consensus guided Network (C3N). Specifically, the “cognitive consensus” refers to the unified-modal label, which conveys the same semantics across the inputs with different modalities. In the AVS field, the “cognitive consensus” means the semantically consistent label of the audio and visual modalities, which is capable of two objectives including estimating the class label of the sounding object and highlighting its pixel-level spatial locations within the frame. By contrast, the traditional audio-visual synchronization or alignment methods [28, 29, 30, 31, 32] aim to match the non-uniform and irregular misaligned audio and visual events from the temporal positions and are unable to provide explicit audio-visual descriptive semantic information. The schematic diagram of our method is shown in Fig 1. The C3N aligns audio and video from the global dimension via semantic-level information and injects the cognitive consensus into local visual regions through an attention mechanism for filling the “dimension gap” and mitigating the mislocalization to improve the segmentation. Firstly, the proposed Cross-modal Cognitive Consensus Inference Module (C3IM) feeds the extracted visual and audio features into independent heads to obtain classification confidence and convert the modality-specific labels into embeddings to calculate the cross-modal semantic similarities. We multiply audio/visual confidence scores and semantic similarities to infer the cognitive consensus and extract a unified-modal label. Next, feed the unified-modal label back to hierarchical layers of the visual backbone as the global semantic-level guidance for highlighting local features corresponding to the object of interest in a channel-spatial attention manner through the parameter-free Cognitive Consensus guided Attention Module (CCAM). Finally, the refined features are fed into a cross-modal feature fusion module and a segmentation head for the segmentation.
@@ II-A Visual Object Segmentation
II-B Sound Source Localization and Segmentation
Recently, researchers attempt to incorporate audio to extract the visual object of interest such as the Sound Source Localization (SSL) task aims to locate the sounding object. SSL evolves from Audio-Visual Correspondence (AVC) [57], which fuses audio and visual features to measure audio-visual consistency. With the growth of deep learning, many SSL methods [58, 59, 60, 61, 62, 63, 64] have been proposed. Senocak et al. [58] adopts contrastive learning for audio-visual feature-level alignment. Chen et al. [62] and Lin et al. [65] improve localization performance through hard negative mining. Some recent methods [60, 61, 63, 64, 66] try to localize in multi-source scenes. Hu et al. [60, 66] propose clustering-based methods. Qian et al. [61] tackle this task by using the class-activated map (CAM). Song et al. [63] propose a negative-free localization method achieved by mining explicit positive examples. Afouras et al. [67], Cheng et al. [68] and Tian et al. [69] propose general audio-visual frameworks, which are suitable for various downstream tasks including SSL. However, SSL generates heat maps for coarsely localizing sounding objects, which lack fine-grained descriptions.
Zhou et al. [14] first proposes the Audio-Visual Segmentation (AVS) task that predicts pixel-level masks for sounding objects, and an AVS baseline utilizes feature-level interactions. Since then, many ideas for AVS have been proposed, such as generation and reconstruction method [24], latent diffusion-based method [23], and representation learning [25]. However, due to the lack of semantic-level guidance, it remains hard to align audio and video explicitly. There are also some methods [26, 27] leverage semantic information in a two-stage strategy (i.e. first generate potential masks, then filter the mask of the sounding object using semantics) and BAVS[27] uses the extra large model to extract semantics, while our C3N is an efficient end-to-end framework and the aligned semantics directly interact with representations of model middle layers to correct and improve the predict segmentation masks.
@@ II-C Label Embedding
In many cases, researchers convert label texts into embedding vectors and combine them with extracted features to enrich the framework’s semantic-level knowledge [75, 76, 77, 78, 79, 80, 81]. Zhang et al. [76] point out the importance of label information and design multiple label embedding-based models. LEAM [77] and EXAM [79] introduce interaction mechanisms to extract input-label relevance clues. GILE [78] proposes a generalized input-label strategy to strengthen the model’s performance on unseen classes. HARNN [82] incorporates recurrent layers with attention mechanism and models multi-layers dependencies. Xiong et al. [80], Cai et al. [83], and Wang et al. [84] introduce label embedding into BERT [74] or its variants to further enhance the semantic-level sensibility of the model. LTTA-LE [85] is a truncation-based method that leverages the label embedding to filter the redundant information of the long text and reduce the text length. In addition, the effectiveness of label embedding is also demonstrated in visual tasks [86, 87] and zero-shot tasks [75].
@@ III Proposed Method
The architecture of our Cross-modal Cognitive Consensus guided Network (C3N) is illustrated in Fig. 2. The network takes an audio and a video as inputs. As described in the pioneering work [14], the audio clips and video frames have been synchronized based on 1-second timestamps by dividing the audio into 1-second clips and re-sampling the frames from the video with 1-second intervals, where is the number of audio signal points, is the length of the audio/video sequence and and are height and width of frames. Then, we use audio/visual encoders to extract their features. The audio features are denoted as , where is the feature dimension. And the visual features are , contains hierarchical features , where . Then, a Cross-modal Cognitive Consensus Inference Module (C3IM) uses pre-trained heads to compute the audio/visual classification confidence and calculate the similarities between audio and visual labels for extracting the semantic-level unified-modal label. Next, the Cognitive Consensus guided Attention Module (CCAM) highlights the local feature elements of the sounding object with the guidance of the inferred unified-modal label. Finally, we feed the refined visual features into the cross-modal feature fusion module and make predictions via the segmentation head.
III-A Cross-modal Cognitive Consensus Inference Module
We aim to explicitly extract global semantic-level alignment information to improve the performance of the AVS framework. Therefore, we propose a novel Cross-modal Cognitive Consensus Inference Module (C3IM) to exploit semantic-level audio/visual classification confidence and labels to infer the audio-visual cognitive consensus. The specific audio-visual semantic alignment mechanism is that the C3IM first conducts modality-specific multi-label classification to extract potential audio and visual elements with their respective classification confidence scores from the input audio and video clips. Second, the audio and visual labels are projected into a unified embedding space to calculate the modality-agnostic label similarities, which facilitates cross-modal semantic-level alignment and overcoming the appearance diversity of objects with identical semantics in complex scenarios. Finally, the C3IM integrates the classification confidence scores and label similarities to obtain the confidence re-weighted matrix, which is used to extract the audio-visual label pair of the sounding object with high semantic consistency and confidence. Since the audio and visual elements are projected into the unified semantic embedding space, the cross-modal alignment of C3IM is not affected by the number of elements.
@@ III-A Cross-modal Cognitive Consensus Inference Module
The visual backbone is initialized with weights pre-trained on ImageNet [88], a large-scale image classification dataset totaling classes. And the audio encoder is pre-trained on AudioSet [89] with audio classes. Such a diversity of classes applies to the semantic description of audio/visual objects in most scenarios. In practice, we freeze the parameters of the audio and visual classification heads. Firstly, the audio features can be represented as , where the superscripts denote timestamps. For the visual feature , where denotes the stage visual feature of the - frame. As shown in Fig. 3, for every-second audio feature and the highest level visual feature , we compute the corresponding classification confidence scores:
@@
where and are pre-trained weights of the audio and visual classification heads, respectively. and denote classification confidence scores, where and are confidence of the audio class and visual class.
@@
Then, to bridge the semantic gap between distinct modalities, we measure the similarities of audio and visual labels and construct a semantic similarity matrix. The AudioSet and ImageNet labels are denoted as and respectively. We convert the audio and visual labels into a unified label embedding space through SpaCy, an advanced library for natural language processing to calculate similarities between label words or phrases, as follows:
where denotes the embedding layer. The audio and visual label embeddings are represented as and respectively, where the embedding dimension is 300. The similarity between the audio and visual labels is defined by the following equation:
@@
where denotes the Frobenius norm, is an element of the semantic similarity matrix , and , denote the row and column indexes.
@@
where and are balance coefficients, which are set to 0.1. Values in the re-weighted matrix can also be considered as cross-modal cognitive consensus degrees. With the inferred cognitive consensus, semantically relevant audio-visual objects can be identified to guide the following segmentation. For ambiguous or noisy audio-visual environments, on the one hand, the unified-modal labels are inferred from the integration and complementarity of the audio and visual modality information, which improves the model’s robustness when a single modality is ambiguous or noisy. On the other hand, previous works [90, 91] have proven that the structures of deep neural networks are able to capture the input data’s statistical prior information and prevent overfitting to the noise.
III-B Cognitive Consensus guided Attention Module
@@
For ease of representation, the time superscript is omitted in this and subsequent sections. After the cross-modal cognitive consensus inference, we feed the semantically consistent unified-modal label to the hierarchical layers of the visual backbone. In addition, we propose a parameter-free Cognitive Consensus guided Attention Module (CCAM), which injects class-activated weights into visual features to highlight the local feature elements corresponding to the sounding object.
We search for the maximum value of and obtain its row and column indexes and :
@@
Specifically, and refer to the index of the audio and visual classes with the highest score. Inspired by the ideology of Grad-CAM [92, 93], we calculate the loss for the visual class and perform the backpropagation to compute the corresponding gradients. Different from Grad-CAM, which computes class-activated heat maps, we aim to obtain the weights corresponding to the unified-modal label with the semantic-level cognitive consensus as follows:
@@
where denotes classification logit of the visual class, denotes feature map activations, denotes the index of the channel, denotes the spatial coordinates of the feature map. In practice, we take the last layer of each stage as the activation layer and the class-activated weights of stage are denoted as , and .
@@
To integrate the class-activated weights with the initial visual features for highlighting the local feature elements corresponding to the cognitive consensus, we propose a parameter-free CCAM, whose structure is shown in the bottom right of Fig 2. Without any learnable parameters, CCAM performs channel-spatial attention to the initial visual features and outputs refined features with the guidance of inferred global semantic-level unified-modal label, as follows:
@@
where denotes the sigmoid function, denotes element-wise multiplication with broadcast mechanism, and cAvg means performing the average operation on the feature along the channel. By means of the channel-spatial attention mechanism, CCAM incorporates semantic-level cognitive consensus weights with feature-level visual representations to guide the model focus on the semantically consistent visual object.
As shown in Fig. 4, although CCAM adopts the well-explored channel-spatial attention mechanism operation similar to the panel (b) of Fig. 4, CCAM utilizes the cognitive consensus inferred by C3IM with explicit semantic information and performs backpropagation to obtain the class-activated weights , and they are then injected into the visual features to highlight the audio-visual semantic consistent object in a channel-spatial attention manner as in the panel (c) of Fig. 4. Unlike the forward propagate feature-level similarity-based cross-modal attention modules [14, 24] as shown in the panel (a) of Fig. 4, our CCAM relies on the attention weights with audio-visual semantic-level alignment information obtained by backpropagation operation.
@@ III-C Cross-modal Feature Fusion
After getting the visual feature , feed it with the audio feature into a cross-modal Non-local block [94, 14], which utilizes 3-D convolutional layers to model the feature-level temporal-spatial dependencies and captures the inter-modal correlations. The cross-modal Non-local block-based feature fusion performs dense feature-level interaction to calibrate the visual features via the discriminative audio features, which enforces highlighting the visual context of sounding objects and ensures accurate segmentation. We first map and repeat into . Then, the cross-modal feature-level interaction process at stage can be formulated as follows:
@@ III-D Segmentation Head
To optimize the parameters of the proposed model. We adopt the BCE loss and Dice loss as the loss function:
@@ IV-A1 Dataset
The proposed method is evaluated on the mainstream AVSbench [14] dataset. For the construction of the AVSBench dataset, Zhou et al. [14] first collect the videos from YouTube using the technique in VGGSound [95] to guarantee the audio-visual intended semantics. Then, depending on the number of sounding objects in the video, the AVSBench dataset is divided into a Single-Source subset and a Multi-Source subset, which are further split into train/val/test sets. Finally, the video frames are annotated with the binary pixel-level masks for the sounding object with 1-second intervals. In practice, each video of AVSBench is cropped to 5 seconds, and video frames at the end of each second are extracted. The Single-Source subset contains 4,392 videos from 23 categories. Note that only the first sampled frame is annotated for videos in the train split, while all sampled frames of the val/test split are annotated, totaling 10,852 annotations. The Multi-Sources subset contains 424 videos and 2,120 annotated pixel-level segmentation masks. All sampled frames of the train/val/test split have annotation since the sounding objects may change over time. As with the pioneering work [14], the following experiments are under the semi-supervised Single Sound Source Segmentation (S4) and the fully supervised Multiple Sound Source Segmentation (MS3) settings.